Two planes start from Los Angeles International Airport and fly in opposite directions. The second plane starts 1/2 hour after the first plane, but its speed is 100 kilometers per hour faster. Find the airspeed of each plane if 2 hours after the first plane departs the planes are 3300 kilometers apart.
Answer by mananth(16949)

plane I speed = x km/h
Plane II speed= x+100
add up speeds
=2x+100
plane I has traveled 0.5x miles before plane II starts
So the distance when plane II starts = (3300-0.5x)
combined speed = 2x+100
time = 3/2 hours (Plane II starts 1/2 hour later. 2-0.5 = 1.5 hours )
3/2 = (3300-0.5x)/(2x+100)
3(2x+100)=2(3300-0.5x)
6x+300=6600-x
7x=6300
x=6300/7
x=900 km/h plane I
Plane II = 1000 km/h
CHECK
plane I distance + plane II distance = 3300
900*2+1000*1.5=3300 km
m.ananth@hotmail.ca